Sparky had room for 16 rows in his garden. He planted 3 rows of tomatoes for every 5 rows of onions. How many rows of tomatoes did he plant?
______rows of tomatoes.
--------------------------------------------------------------------------------
During spring training, Ramons doctor told him that he must drink 1 quart of water for every 3 hours of training he put in. If Ramon drank 3 quarts of water each week during the 6 weeks of training, how many hours did he train?
How many gallons did he drink?
Thanx

Answer by Nate(3500)

Sparky had room for 16 rows in his garden. He planted 3 rows of tomatoes for every 5 rows of onions. How many rows of tomatoes did he plant?
______rows of tomatoes.
Obviously we can see that there are 2 sets of 5 row onions and 2 sets of 3 row tomatoes. There are 6 rows of tomatoes.
--------------------------------------------------------------------------------
During spring training, Ramons doctor told him that he must drink 1 quart of water for every 3 hours of training he put in. If Ramon drank 3 quarts of water each week during the 6 weeks of training, how many hours did he train?
How many gallons did he drink?
3 * 6 = 18 quarts drunk
for 3 quarts, he worked 9 hours, so he worked 54 hours
four quarts in a gallon: 18/4 = 4 2/4 = 4.5 gallons